# Massive deformations of supersymmetric Yang-Mills matrix models

### v1

@@ A.2.2 Type II : $SO(6)\to SO(5)$
$\begin{split}S=\frac{1}{g_{\mathrm{YM}}^{2}}\operatorname{Tr}&\Biggl{[}-\frac{1}{4}[X_{I},X_{J}]^{2}-\frac{i}{2}\epsilon^{ij}\bar{\psi}_{j}\Gamma^{I}[X_{I},\psi_{i}]+2\mu\bar{\psi}_{2}\Gamma^{12345}\psi_{1}+\mu^{2}X_{i}X_{i}+3\mu^{2}X_{6}^{2}\Biggr{]}\,,\end{split}$ (153)
@@
where $M_{ij}$ are $SO(5)$ rotations generators (see (141)) and $T$ is a $U(1)$ generator acting only on the symplectic index of the fermions as
@@
Moreover, the generators $T$ and $M_{ij}$ commute with each other and their commutator with $Q_{a\alpha}$ is non-vanishing. We thus identify the $OSp(2|4)$ superalgebra Freedman:1983na . Since the rotational group in Lorentzian signature becomes either $SO(5)$ or $SO(4,1)$, we write the superalgebra respectively either as $OSp^{*}(2|4)$ or $OSp^{*}(2|2,2)$ following VanProeyen:1999ni .
@@ A.3 Four-dimensional classification
$-m_{0}\epsilon^{IJKL}\tilde{f}_{L}-\frac{1}{2}m_{0}m_{L}\epsilon^{IJKL}-2m_{*}\delta_{[J}^{I}\tilde{f}_{K]}+m_{*}m_{[J}\delta^{I}_{K]}=0\,,\qquad(\Gamma^{JK})$ (168)
@@
We now have two cases. If at least one of the two variables $m_{*}$ and $m_{0}$ is non-vanishing, then everything vanishes except, $m_{0}$, $m_{*}$, $f^{I}_{J}$, $f_{*J}^{I}$ and $S_{IJ}$. We thus have a two-parameter family of solutions $(m,m_{*})$ and the deformation preserves $SO(4)$ symmetry. On the other hand, if both $m_{0}$ and $m_{*}$ vanish, then all these equations impose is that $m^{I}$ and $\tilde{S}^{I}$ point in the same direction. By an $SO(4)$ transformation, we choose $\tilde{S}^{I}=(0,0,0,is)$ and $m^{I}=(0,0,0,m)$. We thus have a two-parameter deformation that breaks $SO(4)\to SO(3)$. Below are written the explicit forms of the solutions.
@@ A.3.1 Type I: $SO(4)\to SO(4)$
$\begin{split}S=\frac{1}{g_{\mathrm{YM}}^{2}}\operatorname{Tr}\Biggl{[}-\frac{1}{4}[X_{I},X_{J}]^{2}-\frac{i}{2}\bar{\psi}\Gamma^{I}[X_{I},\psi]+\mu_{1}\bar{\psi}\psi+i\mu_{2}\bar{\psi}\Gamma_{*}\psi+(\mu_{1}^{2}+\mu_{2}^{2})X_{I}^{2}\Biggr{]}\,,\end{split}$ (171)
@@
$\begin{split}S=\frac{1}{g_{\mathrm{YM}}^{2}}\operatorname{Tr}&\Biggl{[}-\frac{1}{4}[X_{I},X_{J}]^{2}-\frac{i}{2}\bar{\psi}\Gamma^{I}[X_{I},\psi]+\mu\bar{\psi}\psi+\mu^{2}X_{I}^{2}\Biggr{]}\,,\end{split}$ (173)
@@
We identify two separate superalgebras since $Q$ and $Q^{\dagger}$ anti-commute. To understand how the algebra separates, we can define $J_{i}=-\frac{1}{2}\epsilon_{ijk}M_{jk}$, $K_{i}=M_{0i}$, $J^{+}_{i}=J_{i}+iK_{i}$, and $E_{0}=J_{3}^{+}$, $E_{+}=J_{1}^{+}+iJ_{2}^{+}$, $E_{-}=-J_{1}^{+}+iJ_{2}^{+}$. The generators $E_{0}$, $E_{+}$ and $E_{-}$ form a $sl(2,\mathbb{R})$ algebra given by
$[E_{0},E_{\pm}]=\pm E_{\pm}\,,\qquad[E_{+},E_{-}]=-2E_{0}\,.$ (181)
@@ A.3.2 Type II: $SO(4)\to SO(3)$
The second solution breaks $SO(4)\to SO(3)$. Its action reads [^33] Compared to the previous variables, we’ve defined $\mu_{1}\equiv-3\mu m$ and $\mu_{2}\equiv\mu\left(\frac{3s}{2}+3m\right)$
$\begin{split}S=\frac{1}{g_{\mathrm{YM}}^{2}}\operatorname{Tr}&\Biggl{[}-\frac{1}{4}[X_{I},X_{J}]^{2}-\frac{i}{2}\bar{\psi}\Gamma^{I}[X_{I},\psi]-\frac{1}{3}\mu_{1}\bar{\psi}\Gamma^{123}\psi\\ &+i\frac{2}{3}(\mu_{1}+\mu_{2})\epsilon_{ijk}X_{i}X_{j}X_{k}+\left(\frac{2}{9}\mu_{1}^{2}+\frac{1}{3}\mu_{1}\mu_{2}\right)X_{i}X_{i}+\frac{1}{3}\mu_{1}\mu_{2}X_{4}^{2}\Biggr{]}\,,\end{split}$ (183)
and it is invariant the supersymmetry transformations
$\begin{split}\delta_{\epsilon}X^{I}&=\bar{\epsilon}\Gamma^{I}\psi\,,\\ \delta_{\epsilon}\psi&=\frac{i}{2}\Gamma^{IJ}\epsilon[X_{I},X_{J}]+\left(\frac{2}{3}\mu_{1}+\mu_{2}\right)\Gamma^{123}\Gamma^{i}\epsilon X_{i}+\mu_{2}\Gamma^{123}\Gamma^{4}\epsilon X_{4}\,,\end{split}$ (184)
where $i=1,2,3\neq 4$. In order for Myers term to be Hermitian we need $\mu_{1}+\mu_{2}\in\mathbb{R}$ and for the mass terms to be positive we need $\frac{2}{3}\mu_{1}^{2}+\frac{1}{3}\mu_{1}\mu_{2}>0$ and $\mu_{1}\mu_{2}>0$. This leaves two possibilities. The first is $\mu_{1}>0$ and $\mu_{2}>0$ and every term is manifestly fine. The only other possibility is $\mu_{1}=i\mu$ and $\mu_{2}=-i\mu$ which leads to a vanishing Myers term and to the action
@@
This action does not have a real Pfaffian (we would need $\mu\in i\mathbb{R}$ which give an unbounded potential). We thus discard this possibility and restrict to $\mu_{1},\mu_{2}\in\mathbb{R}$ with $\mu_{1}\mu_{2}\geq 0$.
@@ A.4 Three-dimensional classification
$\begin{split}S=\frac{1}{g_{\mathrm{YM}}^{2}}\operatorname{Tr}&\Biggl{[}-\frac{1}{4}[X_{i},X_{j}]^{2}-\frac{i}{2}\bar{\psi}\Gamma^{i}[X_{i},\psi]\\ &+i\mu_{1}\bar{\psi}\psi+\mu_{1}\mu_{2}X_{i}X_{i}+i\frac{2}{3}(\mu_{1}+\mu_{2})\epsilon_{ijk}X_{i}X_{j}X_{k}\Biggr{]}\,,\end{split}$ (193)
@@
where $M_{ij}$ acts as (141) on matrices and $[X_{i},\cdot]$ is a field-dependent gauge transformation. The supercharges transform as spinors under $M_{ij}$. In Lorentzian signature, the $SO(3)$ rotations become $SO(2,1)$. We thus identify the superalgebra $OSp(1|2)$ hardy1979introduction .
@@ Appendix B Pfaffian positivity in four dimensions
It turns out that it is again better to go to Weyl representation, but the action contains terms of the form $\chi^{*}i\sigma^{2}\chi^{*}$ and $-\chi i\sigma^{2}\chi$. Thus the Pfaffian does not reduce to a determinant. Doing so we obtain
@@
Note that since $\mathrm{Pf}(\ldots)^{2}=\mathrm{det(\ldots)}$, we can always write $\mathrm{Pf}(\ldots)=\pm\sqrt{\det(\ldots)}$ locally. Also, it is not hard to see that the Pfaffian is real when $\Omega$ is real. The question is whether $\det(\ldots)$ ever crosses zero. If so, the sign could change. However, by simple manipulations (rearranging columns and changing some signs) one obtains that
$\begin{split}\mathrm{det}\begin{pmatrix}0&2i\Omega&-\mathbf{X}_{3}-i\mathbf{X}_{4}&-\mathbf{X}_{1}+i\mathbf{X}_{2}\\ -2i\Omega&0&-\mathbf{X}_{1}-i\mathbf{X}_{2}&\mathbf{X}_{3}-i\mathbf{X}_{4}\\ -\mathbf{X}_{3}-i\mathbf{X}_{4}&-\mathbf{X}_{1}-i\mathbf{X}_{2}&0&-2i\Omega\\ -\mathbf{X}_{1}+i\mathbf{X}_{2}&\mathbf{X}_{3}-i\mathbf{X}_{4}&2i\Omega&0\end{pmatrix}&=\mathrm{det}\begin{pmatrix}2i\mathrm{1}_{2(N^{2}-1)}&H\\ H^{\dagger}&2i\mathbb{1}_{2(N^{2}-1)}\end{pmatrix}\\ &=\mathrm{det}(H^{\dagger}H+4\Omega^{2}\mathbb{1}_{2(N^{2}-1)})\,,\end{split}$ (201)
@@
This is always non-zero for $\Omega\neq 0$. Thus, the Pfaffian never changes sign when $\Omega\neq 0$. We also know that for $\Omega=0$ it reduces to the Pfaffian of the pure SYM matrix model which we know is positive semi-definite. Thus it is positive semi-definite everywhere and we can write
@@
$\begin{split}\int\left(\prod_{\alpha=1}^{2}\prod_{A=1}^{N^{2}-1}d\chi_{A,\alpha}^{*}d\chi_{A,\alpha}\right)&\mathrm{exp}\left(-\chi_{A}^{\dagger}\left(if_{ABC}X_{I}^{C}\bar{\gamma}^{I}-\frac{2}{3}\Omega_{1}\delta_{AB}\right)\chi_{B}\right)\\ &=\det\underbrace{\left(i\bar{\gamma}^{I}\otimes\mathbf{X}_{I}-\frac{2}{3}\Omega_{1}\mathbb{1}_{2}\otimes\mathbb{1}_{N^{2}-1}\right)}_{\equiv A_{\text{II}}(X)}\,,\end{split}$ (205)
@@ Appendix C Massless limit calculations
Combined with $\Delta(r)$, the $|r_{ab}|$ factors cancel because $01-(D-1)+\mathcal{N}/2=$ 0. Putting all factors together and going back to the original coordinates $r\to\Omega r,\theta\to\sqrt{\Omega}\theta$, we obtain
@@ D.1 Ten-dimensional gamma matrices
where $\gamma^{I=1,\ldots,9}$ are $16\times 16$ real and symmetric matrices satisfying [^35] For an explicit construction, refer to (Freedman:2012zz, , Eq.3.111) or (Komatsu:2024ydh, , App.A) $\{\gamma^{I},\gamma^{J}\}=2\delta^{IJ}$. The $SO(9,1)$ gamma matrices and the chiral matrix $\Gamma_{*}$ are then given by
@@ D.2 Six-dimensional gamma matrices
where $\epsilon^{ij}$ is the Levi-Civita symbol with $\epsilon^{12}=1$. Now this condition is consistent thanks to $\epsilon^{ij}\epsilon^{jk}=-\delta^{ik}$ which effectively cancels the minus sign that we had in $B^{*}B=1$. This condition allows to write $\psi_{1}^{*}$ and $\psi_{2}^{*}$ in terms of $\psi_{1},\psi_{2}$ for a total number of 8 degrees of freedom. This representation of the spinors is the so-called symplectic Majorana-Weyl representation VanProeyen:1999ni ; Kugo:1982bn ; Park:1998nra . Note that it is equivalent to the Weyl representation, since one can use the symplectic condition to write $\psi_{2},\psi_{2}^{*}$ in terms of a single Weyl spinor and its conjugate $\psi_{1},\psi_{1}^{*}$.
@@ Appendix E Matrix integral conventions
Given these generators, one defines the structure constants $f_{ABC}$ from $[T^{A},T^{B}]=if^{ABC}T^{C}$. The measure is defined as
$[dX_{I}]\equiv\prod_{A=1}^{N^{2}}dX_{I}^{A}\,,\qquad[d\psi_{\alpha}]\equiv\prod_{A=1}^{N^{2}}d\psi_{\alpha}^{A}\,,\qquad(U(N))$ (251)
$[dX_{I}]\equiv\prod_{A=1}^{N^{2}-1}dX_{I}^{A}\,,\qquad[d\psi_{\alpha}]\equiv\prod_{A=1}^{N^{2}-1}d\psi_{\alpha}^{A}\,.\qquad(SU(N))$ (252)

### v2

@@ A.2.2 Type II : $SO(6)\to SO(5)$
[…]
@@
where $M_{ij}$ are $SO(5)$ rotation generators (see (141)) and $T$ is a $U(1)$ generator acting only on the symplectic index of the fermions as
@@
Moreover, the generators $T$ and $M_{ij}$ commute with each other and their commutator with $Q_{a\alpha}$ is non-vanishing. We thus identify the $OSp(2|4)$ superalgebra Freedman:1983na. Since the rotational group in Lorentzian signature becomes either $SO(5)$ or $SO(4,1)$, we write the superalgebra as $OSp^{*}(2|4)$ or $OSp^{*}(2|2,2)$, respectively, following VanProeyen:1999ni.
@@ A.3 Four-dimensional classification
$m_{0}\epsilon^{IJKL}\tilde{f}_{L}-\frac{1}{2}m_{0}m_{L}\epsilon^{IJKL}-2m_{*}\delta_{[J}^{I}\tilde{f}_{K]}+m_{*}m_{[J}\delta^{I}_{K]}=0\,,\qquad(\Gamma^{JK})$ (168)
@@
We now have two cases. If at least one of the two variables $m_{*}$ and $m_{0}$ is non-vanishing, then everything vanishes except, $_{0}, mm_{*}$, $f^{I}_{J}$, $f_{*J}^{I}$ and $S_{IJ}$. We thus have a two-parameter family of solutions $(m,m_{*})$ and the deformation preserves $SO(4)$ symmetry. On the other hand, if both $m_{0}$ and $m_{*}$ vanish, then all these equations impose is that $m^{I}$ and $\tilde{S}^{I}$ point in the same direction. By an $SO(4)$ transformation, we choose $\tilde{S}^{I}=(0,0,0,is)$ and $m^{I}=(0,0,0,m)$. We thus have a two-parameter deformation that breaks $SO(4)\to SO(3)$. Below are written the explicit forms of the solutions.
@@ A.3.1 Type I: $SO(4)\to SO(4)$
[…]
@@
[…]
@@
We identify two separate superalgebras since $Q$ and $Q^{\dagger}$ anti-commute. To understand how the algebra separates, we can define $J_{i}=-\frac{1}{2}\epsilon_{ijk}M_{jk}$, $K_{i}=M_{0i}$, $J^{+}_{i}=J_{i}+iK_{i}$, and $E_{0}=J_{3}^{+}$, $E_{+}=J_{1}^{+}+iJ_{2}^{+}$, $E_{-}=-J_{1}^{+}+iJ_{2}^{+}$. The generators $_{0}, E+E_{+}$ and $E_{-}$ form a $sl(2,\mathbb{R})$ algebra given by
$E_{0},E_{\pm}]=\pm E_{\pm}\,,\qquad[E_{+},E_{-}]=-2E_{0}\,.$ (181)
@@ A.3.2 Type II: $SO(4)\to SO(3)$
The second solution breaks $SO(4)\to SO(3)$. Its action reads [^40] Compared to the previous variables, we have defined $\mu_{1}\equiv-3\mu m$ and $\mu_{2}\equiv\mu\left(\frac{3s}{2}+3m\right)$
[…]
and it is invariant under the supersymmetry transformations
[…]
where $i=1,2,3\neq 4$. In order for the Myers term to be Hermitian we need $\mu_{1}+\mu_{2}\in\mathbb{R}$ and for the mass terms to be positive we need $\frac{2}{3}\mu_{1}^{2}+\frac{1}{3}\mu_{1}\mu_{2}>0$ and $\mu_{1}\mu_{2}>0$. This leaves two possibilities. The first is $\mu_{1}>0$ and $\mu_{2}>0$ and every term is manifestly fine. The only other possibility is $\mu_{1}=i\mu$ and $\mu_{2}=-i\mu$ which leads to a vanishing Myers term and to the action
@@
This action does not have a real Pfaffian (we would need $\mu\in i\mathbb{R}$ which would give an unbounded potential). We thus discard this possibility and restrict to $\mu_{1},\mu_{2}\in\mathbb{R}$ with $\mu_{1}\mu_{2}\geq 0$.
@@ A.4 Three-dimensional classification
[…]
@@
where $M_{ij}$ acts as (141) on matrices and $X_{i},\cdot]$ is a field-dependent gauge transformation. The supercharges transform as spinors under $M_{ij}$. In Lorentzian signature, the $SO(3)$ rotations become $SO(2,1)$. We thus identify the superalgebra $OSp(1|2)$ hardy1979introduction.
@@ Appendix B Pfaffian positivity in four dimensions
It turns out that it is again better to go to Weyl representation, but the action contains terms of the form $\chi^{*}i\sigma^{2}\chi^{*}$ and $\chi i\sigma^{2}\chi$. Thus the Pfaffian does not reduce to a determinant. Doing so we obtain
@@
To argue that the Pfaffian is positive, we will use the property $(\mathrm{Pf}\mathcal{M})^{2}=\mathrm{det}\mathcal{M}$. By simple manipulations (rearranging columns and changing some signs) one obtains that
[…]
@@
This is always strictly positive for $\Omega\neq 0$. We thus have
$(\mathrm{Pf}\mathcal{M}_{4,\mathrm{I}}(X))^{2}=\mathrm{det}\mathcal{M}_{4,\mathrm{I}}(X)>0$ (204)
for all $X$ and $\Omega\neq 0$. Since $\mathrm{Pf}\mathcal{M}_{4,\mathrm{I}}(X)$ is a smooth function of $X$ and $\Omega$, and its square is strictly positive, it is real and does not change sign (for all $\Omega^{2}>0$ and all $X$). Also note that at $\Omega=0$ it reduces to the massless case which we already proved to be positive semi-definite. Thus it is positive semi-definite everywhere and we can write
@@
[…]
@@ Appendix C Massless limit calculations
Combined with $\Delta(r)$, the $|r_{ab}|$ factors cancel because $1-(D-1)+\mathcal{N}/2=0$. Putting all factors together and going back to the original coordinates $r\to\Omega r,\theta\to\sqrt{\Omega}\theta$, we obtain
@@ D.1 Ten-dimensional gamma matrices
where $\gamma^{I=1,\ldots,9}$ are $16\times 16$ real and symmetric matrices satisfying [^42] For an explicit construction, refer to (Freedman:2012zz, Eq.3.111) or (Komatsu:2024ydh, App.A) $\{\gamma^{I},\gamma^{J}\}=2\delta^{IJ}$. The $SO(9,1)$ gamma matrices and the chiral matrix $\Gamma_{*}$ are then given by
@@ D.2 Six-dimensional gamma matrices
where $\epsilon^{ij}$ is the Levi-Civita symbol with $\epsilon^{12}=1$. Now this condition is consistent thanks to $\epsilon^{ij}\epsilon^{jk}=-\delta^{ik}$ which effectively cancels the minus sign that we had in $B^{*}B=-1$. This condition allows us to write $\psi_{1}^{*}$ and $\psi_{2}^{*}$ in terms of $\psi_{1},\psi_{2}$ for a total number of 8 degrees of freedom. This representation of the spinors is the so-called symplectic Majorana-Weyl representation VanProeyen:1999ni; Kugo:1982bn; Park:1998nra. Note that it is equivalent to the Weyl representation, since one can use the symplectic condition to write $\psi_{2},\psi_{2}^{*}$ in terms of a single Weyl spinor and its conjugate $\psi_{1},\psi_{1}^{*}$.
@@ Appendix E Matrix integral conventions
Given these generators, one defines the structure constants $f_{ABC}$ from $T^{A},T^{B}]=if^{ABC}T^{C}$. The measure is defined as
$dX_{I}]\equiv\prod_{A=1}^{N^{2}}dX_{I}^{A}\,,\qquad[d\psi_{\alpha}]\equiv\prod_{A=1}^{N^{2}}d\psi_{\alpha}^{A}\,,\qquad(U(N))$ (252)
$dX_{I}]\equiv\prod_{A=1}^{N^{2}-1}dX_{I}^{A}\,,\qquad[d\psi_{\alpha}]\equiv\prod_{A=1}^{N^{2}-1}d\psi_{\alpha}^{A}\,.\qquad(SU(N))$ (253)